OpenStudy (3psilon):

SAT help ? In the equation, K and m are constants . If the equation is true for all values of x, what is the value of m?

OpenStudy (3psilon):

\[(x-8)(x+k) = x^{2} - 5kx +m \]

ganeshie8 (ganeshie8):

hint : 8, -k are the roots of quadratic

OpenStudy (3psilon):

?

hartnn (hartnn):

(x-8)(x+k) multiply this out, what do u get ?

hartnn (hartnn):

(x-8)(x+k)=x^2+(k-8)x-8k = x^2-5kx+m so k-8=5k<------get k from here and put it in m=-8k

hartnn (hartnn):

*k-8=-5k